# Ultra-Faint Milky Way Satellites Discovered in Carina, Phoenix, and Telescopium with DELVE Data Release 3

### v1

@@ 1 INTRODUCTION
Ultra-faint dwarf galaxies (UFDs) are the faintest, oldest, and most dark-matter-dominated stellar systems known. These galaxies have absolute magnitudes fainter than $M_{V}=-7.7$ ($L=10^{5}$ L) (Simon, 2019), and their extremely low luminosity has made them especially powerful probes of the physical processes that regulate galaxy evolution, and the nature of dark matter (e.g., Bullock & Boylan-Kolchin 2017; Simon 2019 and references therein). These systems occupy the extreme limit of galaxy formation and allow us to probe how atomic hydrogen cooling and reionization set the minimum mass of a dark matter halo that is capable of hosting a galaxy (Bullock et al., 2000; Okamoto et al., 2008; Ahvazi et al., 2024; Santos-Santos et al., 2025). On the other hand, the high dark-matter fractions and relatively simple internal dynamics of UFDs let us test dark-matter physics on extremely small spatial scales (e.g., tens of parsecs; Bullock & Boylan-Kolchin 2017). For example, the census of low-mass dark matter subhalos around the Milky Way (as traced by its satellite galaxies) already places lower limits on the dark matter particle mass (Kennedy et al., 2014; Jethwa et al., 2018; Newton et al., 2021; Enzi et al., 2021; Nadler et al., 2020; Newton et al., 2025; Tan et al., 2025a), and UFDs are prime targets for indirect searches for dark matter annihilation or decay (e.g., Ackermann et al., 2015; Geringer-Sameth et al., 2015; McDaniel et al., 2024; Circiello et al., 2025). Nevertheless, because these galaxies are so faint, most of the known population exists as satellites of the Milky Way or of other nearby hosts in the Local Volume (D << 12 Mpc; e.g., Carlsten et al., 2022; Pace, 2025).
@@
The first “ultra-faint” Milky Way satellites were discovered in data from the Sloan Digital Sky Survey (SDSS; e.g., Willman et al., 2005a, b; Belokurov et al., 2007, 2008; Grillmair, 2009). Subsequent wide-field surveys, such as the Pan-STARRS-1 $3\pi$ survey (PS1: Chambers et al., 2016; Laevens et al., 2015a, b), the Dark Energy Survey (DES: DES Collaboration et al., 2016; Bechtol et al., 2015; Koposov et al., 2015a; Kim & Jerjen, 2015a), and the Hyper Suprime-Cam Strategic Survey Program (HSC-SSP; Aihara et al., 2018; Homma et al., 2019, 2024a) pushed the observational frontier to even lower luminosities, increasing the total to $\sim$65 spectroscopically confirmed Milky Way satellite galaxies and likely galaxy candidates (Pace, 2025). The discovery of extremely faint satellites has accelerated due to increasingly sensitive observations; however, many galaxies are still predicted to remain undetected (e.g., Nadler et al., 2020; Tsiane et al., 2025; Tan et al., 2025b).
@@
The faintest confirmed dwarf galaxy, Tucana V, has an absolute magnitude of $M_{V}\sim-1.1$ (Simon et al., 2020; Hansen et al., 2024), corresponding to a stellar mass of $M_{\star}\sim 500$ M. Even fainter and more compact satellites have been reported, including DELVE 1 (Mau et al., 2020), Eridanus III (Simon et al., 2024), and Ursa Major III/UNIONS 1 (Smith et al., 2024) (see Pace 2025 for examples). However, the nature of these systems is unclear because the locus of dwarf galaxies in the size–luminosity plane overlaps that of disrupting star clusters at $M_{V}\gtrsim-3.5$ and velocity dispersions measurements of these faint systems are at the limit of current observational capabilities. Recent theoretical work suggests that extremely low-luminosity galaxies could exist (e.g., Manwadkar & Kravtsov, 2022; Errani et al., 2024a), offering an exciting new regime to test models of galaxy formation and dark matter physics. At the same time, it is likely that some of these systems are star clusters that originated within low-mass galaxies that were subsequently accreted onto the Milky Way. Thus, the nature of ultra-faint compact satellites remains actively debated.
@@
In this paper, we report the discovery of three new ultra-faint satellites in the constellations Carina, Phoenix, and Telescopium. Based on morphological analysis, two of the candidates have physical sizes consistent with the population of Milky Way satellite galaxies. Following established conventions, we designate them Carina IV [^1] After Carina (Cannon et al., 1977), Carina II, and Carina III (Torrealba et al., 2018). and Phoenix III [^2] After Phoenix (Schuster & West, 1976) and Phoenix II (Bechtol et al., 2015; Koposov et al., 2015b).. The third system is more compact and resembles an outer-halo star cluster, thus we follow convention and designate it DELVE 7333After DELVE 1–6 (Mau et al., 2020; Cerny et al., 2021a, 2023a, 2023b)..
This paper is organized as follows. In Section 2 we describe the DELVE data used for our search. In Section 3, we describe our automated search for Milky Way satellite candidates. In Section 4, we characterize the properties of the new systems using data from DELVE and follow-up imaging with Magellan/IMACS for DELVE 7. For the brightest system, Carina IV, we present follow-up observations and analyses that confirm its nature as a Milky Way satellite in Section 5. We briefly discuss the new systems in Section 6 and summarize in Section 7.
2 DELVE Data
The DELVE program (PropID: 2019A-0305) seeks to discover and characterize dwarf galaxy satellites around the Milky Way, Magellanic Clouds, and isolated Magellanic analogs in the Local Group (Drlica-Wagner et al., 2021, 2022). To achieve this, DELVE uses the Dark Energy Camera (DECam; Flaugher et al., 2015) on the Víctor M. Blanco 4-meter Telescope at the NSF Cerro Tololo Inter-American Observatory (CTIO) in Chile to image the high-Galactic-latitude southern sky in the $g$, $r$, $i$, and $z$ bands. To date, DELVE has been allocated more than 150 nights to augment public archival DECam data, including data from DES (DES Collaboration et al., 2016), the DECam Legacy Survey (DECaLS; Dey et al., 2019) and the DECam eROSITA Survey (DeROSITAS; Zenteno et al., 2025). DELVE DR3 (Tan et al. 2025c; Drlica-Wagner et al. in prep.) [^4] https://datalab.noirlab.edu/data/delve includes source catalogs produced from coadded images processed uniformly using the DES Data Management pipeline (DESDM; Morganson et al., 2018), with image detrending and coaddition following DES DR2 (DES Collaboration et al., 2021).
@@ 2 DELVE Data
To identify Milky Way satellites, we search for overdensities consistent with old, metal-poor stellar populations. The DELVE DR3 catalogs include stars belonging to the target Milky Way satellites, foreground Milky Way stars, and distant background galaxies. To obtain a high-purity stellar sample from the catalogs, we select objects with EXT_XGB $\leq 1$. The flag EXT_XGB classifies objects into different morphological classes using the XGBoost algorithm (Chen & Guestrin, 2016; Bechtol et al., 2025b). We also exclude objects with GOLD_FLAGS >0>0 to remove sources with possible data-processing issues or unphysical measurements. Details of both flags can be found in Bechtol et al. (2025b).
@@
Our photometric measurements of the stars come from point-spread function (PSF) model photometry using the PSF_MAG_APER_8 values (Bechtol et al., 2025b). These measurements come from a simultaneous, multi-epoch fit using the individual PSF models for each image at the object location. The flux is normalized to the MAG_APER_8 system to match the convention of SourceExtractor (Bertin & Arnouts, 1996; Bertin et al., 2002). We account for extinction due to interstellar dust by using dereddened measurements with the _CORRECTED suffix, calculated by applying the interstellar extinction correction $A_{b}=R_{b}\times E(B-V)$ where $R_{g}=3.186$, $R_{r}=2.140$, $R_{i}=1.569$, and $R_{z}=1.196$ (DES Collaboration et al., 2021). The E(BV)E(B-V) values are obtained from the Schlegel et al. (1998) reddening maps with the calibration adjustment suggested by Schlafly & Finkbeiner (2011). By convention, we denote extinction-corrected magnitudes with the subscript “0”.
@@
Figure 1: Diagnostic plots for each of the three newly-discovered satellites, closely matching those used to identify the systems in our search. (Top) Stellar density maps for a $1\degree\times 1\degree$ degree region centered on each candidate, smoothed with a $0\farcm 5$ Gaussian kernel. (Middle) Color–magnitude diagrams for stars within $2\farcm 5$ of Carina IV and $1\farcm 5$ of Phoenix III and DELVE 7, centered on each candidate system. Stars are colored by separation from each system’s centroid, with darker colors representing smaller separations. Also overlaid is the best-fit old, metal-poor PARSEC isochrone, with parameters taken from the best-fit ugali results in Table 1. (Bottom) False-color coadded images ($\sim 2\arcmin\times 2\arcmin$, $\sim 0.03\degree\times 0.03\degree$) of Carina IV, Phoenix III, and DELVE 7 from the Legacy Surveys Sky Viewer DR10. Each cutout has been artificially brightened by a factor of $1.8\times$ to highlight faint sources.
3 Dwarf Search Methods
@@
The first search algorithm, simple, has been used to discover at least 29 ultra-faint Milky Way satellites using DECam data (Bechtol et al., 2015; Drlica-Wagner et al., 2015; Mau et al., 2020; Cerny et al., 2021b, a, 2023a, 2023c, 2023b, 2024; Tan et al., 2025c). The simple algorithm first selects stars in color–magnitude space that are consistent with an old ($\tau=12\,\mathrm{Gyr}$), metal-poor ($Z=0.0001$, ${\rm[Fe/H]}\sim-2.2$) PARSEC isochrone (Bressan et al., 2012). This is done by selecting stars that are consistent with $\Delta(g-r)_{0}<\sqrt{0.1^{2}+\sigma_{g}^{2}+\sigma_{r}^{2}},$ where $\sigma_{g}$ and $\sigma_{r}$ are the $g$- and $r$-band magnitude uncertainties of the individual stars, and the constant term $0.1$ accounts for additional modeling and systematic uncertainties. Candidates are then identified by flagging regions that show high densities of isochrone-filtered stars. The simple detection significance of a candidate, SIG, is calculated by comparing the number of observed stars within a circular aperture centered on the candidate, $N_{\rm obs}$, to the surrounding background density field:
@@
The second search algorithm, ugali, employs a maximum-likelihood approach to identify galaxy candidates. Specifically, it compares the likelihood of a model that includes a Milky Way satellite galaxy against a null model consisting only of a uniform field-star background. The likelihood is constructed by evaluating the membership probability of stars within $r<0\fdg 5$ of the candidate system. Each star’s probability of belonging to a putative dwarf galaxy is determined based on its spatial position, photometry, and color. Full details of the likelihood construction are provided in Appendix C of Drlica-Wagner et al. (2020).
@@
Figure 2: The spatial distributions (top) and color–magnitude diagrams (bottom) for likely member stars in Carina IV (left), Phoenix III (center), and DELVE 7 (right). Stars are colored by their membership probabilities derived as part of our ugali characterization analysis, which incorporates the spatial, photometric, and color information of each star. Stars with a lower ugali membership probability ($p<0.05$) are shown in grey. The top panels also show ellipses representing the ugali best-fit half-light radius, $r_{h}$, and three times this half-light radius, $3r_{h}$, for each system. Carina IV is the best-populated of the three systems, featuring a clear main-sequence turn off (MSTO), red giant branch (RGB), and blue horizontal branch (BHB). Phoenix III is at a similar distance, but it is fainter with only a handful of likely member stars. DELVE 7 is substantially fainter, featuring no highly-probable RGB members and is potentially only comprised of main-sequence stars. This lack of RGB stars is consistent with other faint Milky Way satellites at $M_{V}>0$ (e.g., DELVE 5, Kim 3, and Ursa Major III/UNIONS 1).
@@
Due to the large number of hotspots, we initiated a Zooniverse citizen science project “DELVE Dwarf Galaxy Quest: Milky Way Neighbors” to enlist volunteers to visually inspect the hotspots for high-quality galaxy candidates. [^8] https://www.zooniverse.org/projects/ywyh/delve-dwarf-galaxy-quest-milky-way-neighbors Zooniverse is the leading online platform for citizen science and currently has over 1.5 million “citizen scientists” helping to analyze of more than 400 projects (Fortson et al., 2012; Zevin et al., 2017). The Zooniverse volunteers are asked to examine diagnostic plots showing the smoothed stellar density map and color–magnitude diagram of regions around the hotspots, similar to those shown in Figure 1.999The volunteers are not given the Legacy Surveys Sky Viewer images shown in the bottom row of Figure 1. They are then asked to classify each as either a promising candidate for follow-up or a likely false positive. Comprehensive results from this Zooniverse campaign, which will classify over 40,000 hotspots, will be presented in a forthcoming paper.
@@
Here, we present three of the most prominent satellite candidates identified: Carina IV (${\rm SIG_{gr}}=7.5$; $\sqrt{\rm TS_{gr}}=8.8$), Phoenix III (${\rm SIG_{gr}}=4.9$; $\sqrt{\rm TS_{gr}}=6.1$), and DELVE 7 (${\rm SIG_{gr}}=4.9$; $\sqrt{\rm TS_{gr}}=7.5$). Carina IV and DELVE 7 was first identified by DELVE members while assembling the subject set for the Zooniverse project beta test, whereas Phoenix III was discovered by a volunteer beta tester during the same phase. [^11] Phoenix III was first identified by Zooniverse beta tester and citizen scientist Ernest Jude P. Tiu. These candidates generally exhibit much higher simple and ugali detection significances compared to the remaining $\sim$4,000 hotspots detected by both algorithms, placing them within the top 6th percentile in ugali detection significance.
We note that none of the three candidates exceeds the conservative detection thresholds adopted for inclusion in the systematic Milky Way satellite census of Tan et al. (2025b): ${\rm SIG_{gr}}\geq 5.5$, $\sqrt{\rm TS_{gr}}\geq 5.0$, and $\sqrt{\rm TS_{gi}}\geq 5.0$ within the DES footprint, and ${\rm SIG_{gr}}\geq 6.5$, $\sqrt{\rm TS_{gr}}\geq 8.0$, and $\sqrt{\rm TS_{gi}}\geq 7.5$ within the DELVE footprint, where $\sqrt{\rm TS_{gi}}$ is the ugali detection significance in $g$,$i$ bands instead of $g$,$r$ bands. While Carina IV satisfies the census thresholds for $\rm SIG_{gr}$} and $\sqrt{\rm TS_{gr}}$, it falls short of the $\sqrt{\rm TS_{gi}}$ criterion with $\sqrt{\rm TS_{gi}}=6.4$.
4 Characterization of the New Systems
In this section, we present measurements of the structural properties and stellar populations of the three most promising candidates from our search. These measurements are derived primarily with the maximum-likelihood framework of ugali, using the same DECam-based DELVE DR3 data employed in the search (Section 4.1). In addition, Section 4.2 details follow-up Magellan/IMACS imaging obtained for the faintest system, DELVE 7, which provides tighter constraints on its structural parameters and confirms its nature as a metal-poor stellar system.
4.1 Properties from DECam Photometry
We use the maximum-likelihood-based ugali codebase (Bechtol et al., 2015; Drlica-Wagner et al., 2020) to measure the morphological and stellar population properties of the three candidates in the DELVE DR3 data. In the satellite search described in Section 3, we adopted a radially symmetric Plummer profile to reduce computation time, drawing parameters such as size and distance from a predefined grid and fixing a single composite isochrone. Here, we run ugali with a more flexible model that allows both the morphological and stellar population properties to vary freely.

### v2

@@ I INTRODUCTION
Ultra-faint dwarf galaxies (UFDs) are the faintest, oldest, and most dark-matter-dominated stellar systems known. These galaxies have absolute magnitudes fainter than $M_{V}=-7.7$ ($L=10^{5}$ L) (Simon, 2019), and their extremely low luminosity has made them especially powerful probes of the physical processes that regulate galaxy evolution, and the nature of dark matter (e.g., Bullock and Boylan-Kolchin 2017; Simon 2019 and references therein). These systems occupy the extreme limit of galaxy formation and allow us to probe how atomic hydrogen cooling and reionization set the minimum mass of a dark matter halo that is capable of hosting a galaxy (Bullock et al., 2000; Okamoto et al., 2008; Ahvazi et al., 2024; Santos-Santos et al., 2025). On the other hand, the high dark-matter fractions and relatively simple internal dynamics of UFDs let us test dark-matter physics on extremely small spatial scales (e.g., tens of parsecs; Bullock and Boylan-Kolchin 2017). For example, the census of low-mass dark matter subhalos around the Milky Way (as traced by its satellite galaxies) already places lower limits on the dark matter particle mass (Kennedy et al., 2014; Jethwa et al., 2018; Newton et al., 2021; Enzi et al., 2021; Nadler et al., 2020; Newton et al., 2025; Tan et al., 2025c), and UFDs are prime targets for indirect searches for dark matter annihilation or decay (e.g., Ackermann et al., 2015; Geringer-Sameth et al., 2015; McDaniel et al., 2024; Circiello et al., 2025). Nevertheless, because these galaxies are so faint, most of the known population exists as satellites of the Milky Way or of other nearby hosts in the Local Volume (D << 12 Mpc; e.g., Carlsten et al., 2022; Pace, 2025).
@@
The first “ultra-faint” Milky Way satellites were discovered in data from the Sloan Digital Sky Survey (SDSS; e.g., Willman et al., 2005a, b; Belokurov et al., 2007, 2008; Grillmair, 2009). Subsequent wide-field surveys, such as the Pan-STARRS-1 $3\pi$ survey (PS1: Chambers et al., 2016; Laevens et al., 2015b, a), the Dark Energy Survey (DES: DES Collaboration et al., 2016; Bechtol et al., 2015; Koposov et al., 2015; Kim and Jerjen, 2015b), and the Hyper Suprime-Cam Strategic Survey Program (HSC-SSP; Aihara et al., 2018; Homma et al., 2019, 2024) pushed the observational frontier to even lower luminosities, increasing the total to $\sim$65 spectroscopically confirmed Milky Way satellite galaxies and likely galaxy candidates (Pace, 2025). The discovery of extremely faint satellites has accelerated due to increasingly sensitive observations; however, many galaxies are still predicted to remain undetected (e.g., Nadler et al., 2020; Tsiane et al., 2025; Tan et al., 2025b).
@@
The faintest confirmed dwarf galaxy, Tucana V, has an absolute magnitude of $M_{V}\sim-1.1$ (Simon et al., 2020; Hansen et al., 2024), corresponding to a stellar mass of $M_{\star}\sim 500$ M. Even fainter and more compact satellites have been reported, including DELVE 1 (Mau et al., 2020), Eridanus III (Simon et al., 2024), and Ursa Major III/UNIONS 1 (Smith et al., 2024) (see Pace 2025 for examples). However, the nature of these systems is unclear because the locus of dwarf galaxies in the size–luminosity plane overlaps that of disrupting star clusters at $M_{V}\gtrsim-3.5$ and velocity dispersions measurements of these faint systems are at the limit of current observational capabilities. Recent theoretical work suggests that extremely low-luminosity galaxies could exist (e.g., Manwadkar and Kravtsov, 2022; Errani and Peñarrubia, 2020; Errani et al., 2024a; Ahvazi et al., 2025), offering an exciting new regime to test models of galaxy formation and dark matter physics. At the same time, it is likely that some of these systems are star clusters that originated within low-mass galaxies that were subsequently accreted onto the Milky Way. Thus, the nature of ultra-faint compact satellites remains actively debated.
@@
In this paper, we report the discovery of three new ultra-faint satellites in the constellations Carina, Phoenix, and Telescopium. Based on morphological analysis, two of the candidates have physical sizes consistent with the population of Milky Way satellite galaxies. Following established conventions, we designate them Carina IV [^1] After Carina (Cannon et al., 1977), Carina II, and Carina III (Torrealba et al., 2018). and Phoenix III [^2] After Phoenix (Schuster and West, 1976) and Phoenix II (Bechtol et al., 2015; Koposov et al., 2015).. The third system is more compact and resembles an outer-halo star cluster, thus we follow convention and designate it DELVE 7333After DELVE 1–6 (Mau et al., 2020; Cerny et al., 2021a, 2023b, 2023a)..
This paper is organized as follows. In Section II we describe the DELVE data used for our search. In Section III, we describe our automated search for Milky Way satellite candidates. In Section IV, we characterize the properties of the new systems using data from DELVE and follow-up imaging with Magellan/IMACS for DELVE 7. For the brightest system, Carina IV, we present follow-up observations and analyses that confirm its nature as a Milky Way satellite in Section V. We briefly discuss the new systems in Section VI and summarize in Section VII.
II DELVE Data
The DELVE program (PropID: 2019A-0305) seeks to discover and characterize dwarf galaxy satellites around the Milky Way, Magellanic Clouds, and isolated Magellanic analogs in the Local Group (Drlica-Wagner et al., 2021, 2022). To achieve this, DELVE uses the Dark Energy Camera (DECam; Flaugher et al., 2015) on the Víctor M. Blanco 4-meter Telescope at the NSF Cerro Tololo Inter-American Observatory (CTIO) in Chile to image the high-Galactic-latitude southern sky in the $g$, $r$, $i$, and $z$ bands. To date, DELVE has been allocated more than 150 nights to augment public archival DECam data, including data from DES (DES Collaboration et al., 2016), the DECam Legacy Survey (DECaLS; Dey et al., 2019) and the DECam eROSITA Survey (DeROSITAS; Zenteno et al., 2025). DELVE DR3 (Tan et al. 2025a; Drlica-Wagner et al. in prep.) [^4] https://datalab.noirlab.edu/data/delve includes source catalogs produced from coadded images processed uniformly using the DES Data Management pipeline (DESDM; Morganson et al., 2018), with image detrending and coaddition following DES DR2 (DES Collaboration et al., 2021).
@@ II DELVE Data
To identify Milky Way satellites, we search for overdensities consistent with old, metal-poor stellar populations. The DELVE DR3 catalogs include stars belonging to the target Milky Way satellites, foreground Milky Way stars, and distant background galaxies. To obtain a high-purity stellar sample from the catalogs, we select objects with EXT_XGB $\leq 1$. The flag EXT_XGB classifies objects into different morphological classes using the XGBoost algorithm (Chen and Guestrin, 2016; Bechtol et al., 2025), with integer classes ranging from 0 (high-confidence stars) to 4 (high-confidence galaxies). As an additional check, we examine galaxy density maps around the candidates, constructed using non-stellar sources with $\texttt{EXT\_XGB}>2$, to ensure that the detected overdensities arise from genuine stellar systems rather than from misclassified galaxies. We also exclude objects with GOLD_FLAGS >0>0 to remove sources with possible data-processing issues or unphysical measurements. Details of both flags can be found in Bechtol et al. (2025).
@@
Our photometric measurements of the stars come from point-spread function (PSF) model photometry using the PSF_MAG_APER_8 values (Bechtol et al., 2025). These measurements come from a simultaneous, multi-epoch fit using the individual PSF models for each image at the object location. The flux is normalized to the MAG_APER_8 system to match the convention of SourceExtractor (Bertin and Arnouts, 1996; Bertin et al., 2002). We account for extinction due to interstellar dust by using dereddened measurements with the _CORRECTED suffix, calculated by applying the interstellar extinction correction $A_{b}=R_{b}\times E(B-V)$ where $R_{g}=3.186$, $R_{r}=2.140$, $R_{i}=1.569$, and $R_{z}=1.196$ (DES Collaboration et al., 2021). The E(BV)E(B-V) values are obtained from the Schlegel et al. (1998) reddening maps with the calibration adjustment suggested by Schlafly and Finkbeiner (2011). By convention, we denote extinction-corrected magnitudes with the subscript “0”.
@@
Figure 1: Diagnostic plots for each of the three newly-discovered satellites, closely matching those used to identify the systems in our search. (Top) Isochrone filtered stellar density maps for a $1\degree\times 1\degree$ degree region centered on each candidate, smoothed with a $0\farcm 5$ Gaussian kernel. (Middle) Color–magnitude diagrams for stars within $2\farcm 5$ of Carina IV and $1\farcm 5$ of Phoenix III and DELVE 7, centered on each candidate system. Stars are colored by separation from each system’s centroid, with darker colors representing smaller separations. Also overlaid is the best-fit old, metal-poor PARSEC isochrone, with parameters taken from the best-fit ugali results in Table 1. The median photometric uncertainties for stars around each satellite, as a function of $g$-band magnitude, are indicated by red error bars. The “x” marker indicates a confirmed non–member star of DELVE 7 with a Gaia parallax of $\varpi=0.71\pm 0.15$ mas, clearly placing it in the foreground of the system. (Bottom) False-color coadded images ($\sim 2\arcmin\times 2\arcmin$, $\sim 0.03\degree\times 0.03\degree$) of Carina IV, Phoenix III, and DELVE 7 from the Legacy Surveys Sky Viewer DR10. Each cutout has been artificially brightened by a factor of $1.8\times$ to highlight faint sources. We also overlay red ellipses on the coadded images demarcating the half-light radius, $r_{h}$, of the satellites.
III Dwarf Search Methods
@@
The first search algorithm, simple, has been used to discover at least 29 ultra-faint Milky Way satellites using DECam data (Bechtol et al., 2015; Drlica-Wagner et al., 2015; Mau et al., 2020; Cerny et al., 2021b, a, 2023b, 2023c, 2023a, 2024; Tan et al., 2025a). The simple algorithm first selects stars in color–magnitude space that are consistent with an old ($\tau=12\,\mathrm{Gyr}$), metal-poor ($Z=0.0001$, ${\rm[Fe/H]}\sim-2.2$) PARSEC isochrone (Bressan et al., 2012). This is done by selecting stars that are consistent with $\Delta(g-r)_{0}<\sqrt{0.1^{2}+\sigma_{g}^{2}+\sigma_{r}^{2}},$ where $\sigma_{g}$ and $\sigma_{r}$ are the $g$- and $r$-band magnitude uncertainties of the individual stars, and the constant term $0.1$ accounts for additional modeling and systematic uncertainties. Candidates are then identified by flagging regions that show high densities of isochrone-filtered stars. The simple detection significance of a candidate, SIG, is calculated by comparing the number of observed stars within a circular aperture centered on the candidate, $N_{\rm obs}$, to the surrounding background density field:
@@
The second search algorithm, ugali, employs a maximum-likelihood approach to identify galaxy candidates. Specifically, it compares the likelihood of a model that includes a Milky Way satellite galaxy against a null model consisting only of a uniform field-star background. The uniform field-star background population is determined empirically using stars from a circular annulus surrounding each candidate ($0\fdg 5<r<2\fdg 0$), where they are binned in color–magnitude space. The likelihood is constructed by evaluating the membership probability of stars within $r<0\fdg 5$ of the candidate system. Each star’s probability of belonging to a putative dwarf galaxy is determined based on its spatial position, photometry, and color. Full details of the likelihood construction are provided in Bechtol et al. (2015) and Appendix C of Drlica-Wagner et al. (2020).
@@
Figure 2: The spatial distributions (top) and color–magnitude diagrams (bottom) for likely member stars in Carina IV (left), Phoenix III (center), and DELVE 7 (right). Stars are colored by their membership probabilities derived as part of our ugali characterization analysis, which incorporates the spatial, photometric, and color information of each star. Stars with a lower ugali membership probability ($p_{\rm ugali}<0.05$) are shown in grey. The top panels also show ellipses representing the ugali best-fit half-light radius, $r_{h}$, and three times this half-light radius, $3r_{h}$, for each system. For Carina IV, we highlight with red outlines the stars in the color–magnitude diagram that have both a high ugali membership probability ($p_{\rm ugali}>0.20$) and a Gaia detection. These stars are discussed further in Section V.1 and Appendix A. Carina IV is the best-populated of the three systems, featuring a clear main-sequence turn off (MSTO), red giant branch (RGB), and blue horizontal branch (BHB). Phoenix III is at a similar distance, but it is fainter with only a handful of likely member stars. DELVE 7 is substantially fainter, featuring no highly-probable RGB members and is potentially only comprised of main-sequence stars. This lack of RGB stars is consistent with other faint Milky Way satellites at $M_{V}>0$ (e.g., DELVE 5, Kim 3, and Ursa Major III/UNIONS 1).
@@
Due to the large number of hotspots, we initiated a Zooniverse citizen science project “DELVE Dwarf Galaxy Quest: Milky Way Neighbors” to enlist volunteers to visually inspect the hotspots for high-quality galaxy candidates. [^8] https://www.zooniverse.org/projects/ywyh/delve-dwarf-galaxy-quest-milky-way-neighbors Zooniverse is the leading online platform for citizen science and currently has over 1.5 million “citizen scientists” helping to analyze more than 400 projects (Fortson et al., 2012; Zevin et al., 2017). The Zooniverse volunteers are asked to examine diagnostic plots showing the smoothed stellar density map and color–magnitude diagram of regions around the hotspots, similar to those shown in Figure 1.999The volunteers are not given the Legacy Surveys Sky Viewer images shown in the bottom row of Figure 1. They are then asked to classify each as either a promising candidate for follow-up or a likely false positive. Comprehensive results from this Zooniverse campaign, which will classify over 40,000 hotspots, will be presented in a forthcoming paper.
@@
Here, we present three of the most prominent satellite candidates identified: Carina IV (${\rm SIG_{gr}}=7.5$; $\sqrt{\rm TS_{gr}}=8.8$), Phoenix III (${\rm SIG_{gr}}=4.9$; $\sqrt{\rm TS_{gr}}=6.1$), and DELVE 7 (${\rm SIG_{gr}}=4.9$; $\sqrt{\rm TS_{gr}}=7.5$). Carina IV and DELVE 7 were first identified by DELVE members, based on the presence of clear Milky Way satellite signatures in their diagnostic plots (Fig. 1), including a prominent overdensity in the stellar density map and a well-defined color–magnitude sequence consistent with a metal-poor stellar population. In contrast, Phoenix III was discovered by a volunteer beta tester when going through the same diagnostic plots during the same phase. [^11] Phoenix III was first identified by Zooniverse beta tester and citizen scientist Ernest Jude P. Tiu. These candidates generally exhibit much higher simple and ugali detection significances compared to the remaining $\sim$4,000 hotspots detected by both algorithms, placing them within the top 6th percentile in ugali detection significance.
We note that none of the three candidates exceeds the conservative detection thresholds adopted for inclusion in the systematic Milky Way satellite census of Tan et al. (2025b). [^12] The detection threshold is ${\rm SIG_{gr}}\geq 5.5$, $\sqrt{\rm TS_{gr}}\geq 5.0$, and $\sqrt{\rm TS_{gi}}\geq 5.0$ within the DES footprint, and ${\rm SIG_{gr}}\geq 6.5$, $\sqrt{\rm TS_{gr}}\geq 8.0$, and $\sqrt{\rm TS_{gi}}\geq 7.5$ within the DELVE footprint, where $\sqrt{\rm TS_{gi}}$ is the ugali detection significance in $g$,$i$ bands instead of $g$,$r$ bands. While Carina IV satisfies the census thresholds for $\rm SIG_{gr}$} and $\sqrt{\rm TS_{gr}}$, it falls short of the $\sqrt{\rm TS_{gi}}$ criterion with $\sqrt{\rm TS_{gi}}=6.4$. In contrast to Tan et al. (2025b), which prioritizes high dwarf galaxy sample purity through the use of additional $i$ band information, the search and the main analysis presented in this work rely solely on $g$- and $r$-band data.
IV Characterization of the New Systems
In this section, we present measurements of the structural properties and stellar populations of the three most promising candidates from our search. These measurements are derived primarily with the maximum-likelihood framework of ugali, using the same DECam-based DELVE DR3 data employed in the search (Section IV.1). In addition, Section IV.2 details follow-up Magellan/IMACS imaging obtained for the faintest system, DELVE 7, which provides tighter constraints on its structural parameters and confirms its nature as a metal-poor stellar system.
IV.1 Properties from DECam Photometry
We use the maximum-likelihood-based ugali codebase (Bechtol et al., 2015; Drlica-Wagner et al., 2020) to measure the morphological and stellar population properties of the three candidates in the DELVE DR3 data. In the satellite search described in Section III, we adopted a radially symmetric Plummer profile to reduce computation time, drawing parameters such as size and distance from a predefined grid and fixing a single composite isochrone. Here, we run ugali with a more flexible model that allows both the morphological and stellar population properties to vary freely.